Assume that when the earth was created there was the same amount of 235 U and 238 U; the present ratio of 235 U to 238 U is 0.7%. The half life of 235 U is 8.8x10^8 yr. and for 238 U is 4.5 x 10^9 yr. What is the age of the Earth?
Answer #1

Remaining Amount o - t/th N= No (2) -t/til N = Noz (2) e1th N Noz (2) - t/t1/ U235 6235 = 0.7 → 238 x 100 = = 7x10-3 7x10-3 u